# Ill-posedness issue on the Oldroyd-B model in the critical Besov spaces

### v1

@@ 3.1 Construction of initial data
    Then we let
@@
    $\left(A\right)_{ij}:=\begin{cases}\varepsilon,\quad 1\leq i=j\leq 2,\\ 1,\quad 3\leq i=j\leq d,\\ 0,\quad else.\end{cases}$
@@
$\widehat{b_{n}}(\xi)=$ $\frac{\mathrm{i}\Gamma_{n}2^{n}}{\varepsilon^{2}2^{dn+1}}\left\{\prod_{i=1}^{2}\widehat{\theta}\left(\frac{\xi_{i}+\widetilde{\lambda}_{n}}{\varepsilon 2^{n}}\right)-\prod_{i=1}^{2}\widehat{\theta}\left(\frac{\xi_{i}-\widetilde{\lambda}_{n}}{\varepsilon 2^{n}}\right)\right\}\prod_{i=3}^{d}\widehat{\theta}\left(\frac{\xi_{i}}{2^{n}}\right),$
@@
Obviously, from (3.8), one has $\mathrm{div}\ u^{n}_{0}=0$.
@@ 3.2 Estimation of initial data
[…]
@@
$\|\dot{\Delta}_{n}\left(\tau^{n}_{0}\omega^{n}_{0}-\omega^{n}_{0}\tau^{n}_{0}\right)\|_{L^{\infty}(\mathbb{R}^{d})}=\|\tau^{n}_{0}\omega^{n}_{0}-\omega^{n}_{0}\tau^{n}_{0}\|_{L^{\infty}(\mathbb{R}^{d})}\geq\tilde{c}2^{2n}\Gamma^{2}_{n}\ln n.$
@@
and the support condition of $b_{n}$ and $d_{n}$, then
@@
and
$\|\partial_{1}c_{n}d_{n}\|_{L^{\infty}}\leq\|\partial_{1}c_{n}\|_{L^{\infty}}\|d_{n}\|_{L^{\infty}}\leq\|\xi_{1}\widehat{c_{n}}(\xi)\|_{L^{1}}\|d_{n}\|_{L^{\infty}}\leq\tilde{C}\Gamma^{2}_{n}\varepsilon 2^{2n}\sum^{n/2}_{j=1}\frac{1}{j},$ (3.14)
where we have used the fact $|\xi_{1}-\xi_{2}|\leq\tilde{C}\varepsilon 2^{n}$ and $|\xi_{1}|\approx|\xi_{2}|\approx 2^{n}$ when $\xi\in\mathrm{supp}\ \widehat{c_{n}}$. Here the constants $\tilde{c}$ and $\tilde{C}$ do not depend on the parameter $\varepsilon$.
@@
Let $(u^{n}_{0},\tau^{n}_{0})$ be defined by (3.9)-(3.10). Given $1\leq p<2d$, there exist some constant $C_{0}>1$ which may depend on $\varepsilon$ and $\phi$ but not $n$, and $N_{0}$ such that for $n>N_{0}$, System (1.1) has a unique local solution $(u,\tau)$ associated with initial data $(u^{n}_{0},\tau^{n}_{0})$ satisfying
$u\in\mathcal{C}\left([0,T],\dot{B}_{p,1}^{\frac{d}{p}-1}\cap\dot{B}_{p,1}^{\frac{d}{p}+3}\right)\cap{L}^{1}\left([0,T],\dot{B}_{p,1}^{\frac{d}{p}+1}\cap\dot{B}_{p,1}^{\frac{d}{p}+5}\right),$
$\tau\in\mathcal{C}\left([0,T],\dot{B}_{p,1}^{\frac{d}{p}}\cap\dot{B}_{p,1}^{\frac{d}{p}+4}\right),$
@@
$\|u\|_{L^{\infty}_{T}(\dot{B}^{\frac{d}{p}-1+k}_{p,1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{p}+1+k}_{p,1})}\leq C_{0}\Gamma_{n}2^{kn},$
$\|\tau\|_{L^{\infty}_{T}(\dot{B}^{\frac{d}{p}+k}_{{p},1})}\leq C_{0}\Gamma_{n}\ln n2^{kn}.$
@@
$u\in\mathcal{C}\left([0,T],\dot{B}_{p,1}^{\frac{d}{p}-1}\right)\cap{L}^{1}\left([0,T],\dot{B}_{p,1}^{\frac{d}{p}+1}\right),\quad\tau\in\mathcal{C}\left([0,T],\dot{B}_{p,1}^{\frac{d}{p}}\right).$
@@
$\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}-1}_{{p},1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{p}+1}_{p,1})}\leq C\left(\|u^{n}_{0}\|_{\dot{B}^{\frac{d}{p}-1}_{p,1}}+\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}-1}_{p,1})}\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{p}+1}_{{p},1})}+T\|\tau\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}}_{{p},1})}\right).$ (3.15)
@@
$\|\tau\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}}_{{p},1})}\leq\exp\left(C\|u\|_{L_{T}^{1}(\dot{B}_{{p},1}^{\frac{d}{{p}}+1})}\right)\left(\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{{p},1}}+C\|\tau\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}}_{{p},1})}\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{p}+1}_{{p},1})}\right).$ (3.16)
@@
$X_{T}=\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}-1}_{{p},1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{p}+1}_{{p},1})},\quad Y_{T}=\|\tau(t,\cdot)\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}}_{{p},1})}\quad\text{and}\quad Z_{T}:=X_{T}+2CTY_{T}.$
@@
$Z_{T}$ $\leq\left(C+Ce^{CX_{T}}\right)\left(\|u^{n}_{0}\|_{\dot{B}^{\frac{d}{p}-1}_{{p},1}}+T\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{{p},1}}+(Z_{T})^{2}\right)\leq Ce^{CZ_{T}}\left(\Gamma_{n}+(Z_{T})^{2}\right).$ (3.17)
@@
$\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}-1}_{{p},1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{p}+1}_{{p},1})}\leq Z_{T}\leq 2C\Gamma_{n}.$ (3.18)
@@
$\|\tau\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}}_{{p},1})}\leq 2\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{{p},1}}\leq C\Gamma_{n}\ln n.$ (3.19)
@@
$\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{{p},1}^{\frac{d}{p}+1})}$ $\leq\exp\left(C\|u\|_{L_{T}^{1}(\dot{B}_{{p},1}^{\frac{d}{{p}}+1})}\right)\left(\left\|\tau^{n}_{0}\right\|_{\dot{B}_{{p},1}^{\frac{d}{p}+1}}+C\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{p,1}^{\frac{d}{p}})}\|u\|_{L_{T}^{1}(\dot{B}_{{p},1}^{\frac{d}{{p}}+2})}\right.$
$\quad\left.+C\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{p,1}^{\frac{d}{p}+1})}\|u\|_{L_{T}^{1}(\dot{B}_{p,1}^{\frac{d}{p}+1})}\right)$
$\leq C\Gamma_{n}\left(2^{n}\ln n+\ln n\|u\|_{\tilde{L}_{T}^{1}(\dot{B}_{{p},1}^{\frac{d}{{p}}+2})}+\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{p,1}^{\frac{d}{p}+1})}\right),$
@@
$\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{{p},1}^{\frac{d}{p}+1})}$ $\leq C\Gamma_{n}\ln n\left(2^{n}+\|u\|_{\tilde{L}_{T}^{1}(\dot{B}_{{p},1}^{\frac{d}{{p}}+2})}\right).$ (3.20)
@@
$\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}}_{{p},1})}$ $+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{p}+2}_{{p},1})}\leq\left\|u^{n}_{0}\right\|_{\dot{B}_{p,1}^{\frac{d}{p}}}+C\|u\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{p,1}^{\frac{d}{p}})}\|u\|_{L_{T}^{1}(\dot{B}_{p,1}^{\frac{d}{p}+1})}+T\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{p,1}^{\frac{d}{p}+1})}$
$\leq C\left\{\Gamma_{n}2^{n}+\Gamma_{n}^{-1}2^{-n}+(\Gamma_{n}+\Gamma_{n}^{-1}2^{-2n})\left(\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}}_{{p},1})}+\|u\|_{\tilde{L}_{T}^{1}(\dot{B}_{{p},1}^{\frac{d}{{p}}+2})}\right)\right\},$
@@
$\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{p}}_{{p},1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{p}+2}_{{p},1})}$ $\leq C\Gamma_{n}2^{n}.$
@@
$\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{{p},1}^{\frac{d}{p}+1})}$ $\leq C\Gamma_{n}\ln n2^{n}.$
@@ 3.4 Discontinuity of the solution map
Now, we give the lower bound estimation of $\|\tau(T)\|_{\dot{B}^{\frac{d}{p}}_{p,r}}$ with $T=2^{-2n}(\ln n\Gamma_{n}^{2})^{-1}$ which is crucial for the proof of the discontinuity of solutions.
@@
$\|\tau(t)-\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{{p},1}}$ $\leq Ct\|\tau\|_{\tilde{L}_{t}^{\infty}\dot{B}^{\frac{d}{p}}_{{p},1}}\|u\|_{\tilde{L}_{t}^{\infty}\dot{B}^{\frac{d}{p}+1}_{{p},1}}+Ct\|\tau\|_{\tilde{L}_{t}^{\infty}\dot{B}^{\frac{d}{p}+1}_{{p},1}}\|u\|_{\tilde{L}_{t}^{\infty}\dot{B}^{\frac{d}{p}}_{{p},1}}$
@@
$\|\omega(t)-\omega^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{{p},1}}$ $\leq C\|u(t)-u^{n}_{0}\|_{\dot{B}^{\frac{d}{p}+1}_{{p},1}}$
$\leq C\|u\|_{\tilde{L}_{t}^{\infty}\dot{B}^{\frac{d}{p}}_{{p},1}}\|u\|_{\tilde{L}_{t}^{1}\dot{B}^{\frac{d}{p}+2}_{{p},1}}+Ct\|u\|_{\tilde{L}_{t}^{\infty}\dot{B}^{\frac{d}{p}+3}_{p,1}}+Ct\|\tau\|_{\tilde{L}_{t}^{\infty}\dot{B}^{\frac{d}{p}+2}_{{p},1}}$
$\leq C\Gamma^{2}_{n}2^{2n}+Ct\Gamma_{n}2^{4n}+Ct\Gamma_{n}\ln n2^{2n}\leq C\Gamma^{2}_{n}2^{2n}+C(\ln n\Gamma_{n})^{-1}2^{2n},$
@@
$\|P(t)-P_{0}\|_{L^{\infty}}$ $\leq C\|\omega(t)-\omega^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{{p},1}}\|\tau(t)\|_{\tilde{L}_{t}^{\infty}(\dot{B}_{p,1}^{\frac{d}{p}})}+C\|\tau(t)-\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{p,1}}\|u\|_{\tilde{L}_{t}^{\infty}(\dot{B}_{p,1}^{\frac{d}{p}+1})}$
@@
$\leq$ $~{}CT^{2}\left(\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}+1}_{p,1}}\|u^{n}_{0}\|_{\dot{B}^{\frac{d}{p}+1}_{p,1}}+\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{p,1}}\|u^{n}_{0}\|_{\dot{B}^{\frac{d}{p}+2}_{p,1}}\right)\|u\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{p,1}^{\frac{d}{p}})}$
@@
$T\|P_{0}\|_{\dot{B}^{\frac{d}{p}}_{p,r}}\geq\tilde{c}T2^{2n}\ln n\Gamma_{n}^{2}\quad\text{and}\quad\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{p,r}}\leq C\Gamma_{n}.$ (3.23)

### v2

@@ 3.1 Construction of initial data
    We assume that $\widehat{\theta}$ is even and real-valued such that $\theta$ is a real-valued function. Then we can define the new real-valued function
@@
    $d=2,\quad\left(A\right)_{ij}:=\begin{cases}\varepsilon,\quad 1\leq i=j\leq d,\\ 0,\quad else,\end{cases}$
    $d\geq 3,\quad\left(A\right)_{ij}:=\begin{cases}\varepsilon,\quad 1\leq i=j\leq 2,\\ 1,\quad 3\leq i=j\leq d,\\ 0,\quad else.\end{cases}$
@@
$\widehat{b_{n}}(\xi)=$ $\frac{\mathrm{i}\Gamma_{n}2^{n}}{\varepsilon^{2}2^{dn+1}}\left\{\prod_{i=1}^{2}\widehat{\theta}\left(\frac{\xi_{i}+\widetilde{\lambda}_{n}}{\varepsilon 2^{n}}\right)-\prod_{i=1}^{2}\widehat{\theta}\left(\frac{\xi_{i}-\widetilde{\lambda}_{n}}{\varepsilon 2^{n}}\right)\right\}\prod_{i=3}^{d}\widehat{\theta}\left(\frac{\xi_{i}}{2^{n}}\right),$ (3.9)
@@
From (3.9), we also have
$\mathrm{supp}\ \widehat{b_{n}}(\xi)\subset\bf{S}_{1}\cup\bf{S}_{2},$ (3.11)
where
[…]
@@
Obviously, from (3.10), one has $\mathrm{div}\ u^{n}_{0}=0$. We would like to emphasize that both initial data $u^{n}_{0}$ and $\tau^{n}_{0}$ are real-valued Schwarz functions. Next, we need to verify that for large $n$ enough
  • initial data $u^{n}_{0}$ and $\tau^{n}_{0}$ is small in $\dot{B}^{{d}/{p}-1}_{p,r}\times\dot{B}^{{d}/{p}}_{p,r}$ for any $r>1$;
  • the main contributor $\omega^{n}_{0}\tau^{n}_{0}-\tau^{n}_{0}\omega^{n}_{0}$ is large in $L^{\infty}$.
@@ 3.2 Estimation of initial data
$\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{p,1}}\leq C\Gamma_{n}\ln n\quad\text{and}\quad\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{p,r}}\leq C\Gamma_{n}\quad\text{for}\;1<r\leq\infty.$
@@
$\left\|\dot{\Delta}_{n}\left(\tau^{n}_{0}\omega^{n}_{0}-\omega^{n}_{0}\tau^{n}_{0}\right)\right\|_{L^{\infty}(\mathbb{R}^{d})}=\|\tau^{n}_{0}\omega^{n}_{0}-\omega^{n}_{0}\tau^{n}_{0}\|_{L^{\infty}(\mathbb{R}^{d})}\geq\tilde{c}2^{2n}\Gamma^{2}_{n}\ln n.$
@@
and the support conditions of $\widehat{b_{n}}$ and $\widehat{d_{n}}$, then
@@
Recalling the construction of $(b_{n},c_{n},d_{n})$ and (3.9), we deduce that
$\|\partial_{1}c_{n}d_{n}\|_{L^{\infty}}$ $\leq\|\partial_{1}c_{n}\|_{L^{\infty}}\|d_{n}\|_{L^{\infty}}\leq\|\xi_{1}\widehat{c_{n}}(\xi)\|_{L^{1}}\|d_{n}\|_{L^{\infty}}$
$\leq\left\|\frac{\xi_{1}(\xi_{1}-\xi_{2})}{\xi_{2}}\widehat{b_{n}}\right\|_{L^{1}}\|d_{n}\|_{L^{\infty}}$
$\leq\tilde{C}\varepsilon 2^{n}\|\widehat{b_{n}}\|_{L^{1}}\|d_{n}\|_{L^{\infty}}\leq\tilde{C}\Gamma^{2}_{n}\varepsilon 2^{2n}\sum^{n/2}_{j=1}\frac{1}{j},$ (3.17)
where we have used (3.11). In fact, (3.11) implies that $|\xi_{1}-\xi_{2}|\leq\tilde{C}\varepsilon 2^{n}$ and $|\xi_{1}|\approx|\xi_{2}|\approx 2^{n}$ when $\xi\in\mathrm{supp}\ \widehat{b_{n}}$. Here we would like to emphasize that the above constants $\tilde{c}$ and $\tilde{C}$ do not depend on the parameter $\varepsilon$.
@@
Let $(u^{n}_{0},\tau^{n}_{0})$ be defined by (3.12)-(3.13). Given $1\leq q<2d$, there exist some constant $C_{0}>1$ which may depend on $\varepsilon$ and $\phi$ but not $n$, and $N_{0}$ such that for $n>N_{0}$, System (1.1) has a unique local solution $(u,\tau)$ associated with initial data $(u^{n}_{0},\tau^{n}_{0})$ satisfying
$u\in\mathcal{C}\left([0,T],\dot{B}_{q,1}^{\frac{d}{q}-1}\cap\dot{B}_{q,1}^{\frac{d}{q}+3}\right)\cap{L}^{1}\left([0,T],\dot{B}_{q,1}^{\frac{d}{q}+1}\cap\dot{B}_{q,1}^{\frac{d}{q}+5}\right),$
$\tau\in\mathcal{C}\left([0,T],\dot{B}_{q,1}^{\frac{d}{q}}\cap\dot{B}_{q,1}^{\frac{d}{q}+4}\right),$
@@
$\|u\|_{L^{\infty}_{T}(\dot{B}^{\frac{d}{q}-1+k}_{q,1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{q}+1+k}_{q,1})}\leq C_{0}\Gamma_{n}2^{kn},$
$\|\tau\|_{L^{\infty}_{T}(\dot{B}^{\frac{d}{q}+k}_{{q},1})}\leq C_{0}\Gamma_{n}\ln n2^{kn}.$
@@
$u\in\mathcal{C}\left([0,T],\dot{B}_{q,1}^{\frac{d}{q}-1}\right)\cap{L}^{1}\left([0,T],\dot{B}_{q,1}^{\frac{d}{q}+1}\right),\quad\tau\in\mathcal{C}\left([0,T],\dot{B}_{q,1}^{\frac{d}{q}}\right).$
@@
$\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}-1}_{{q},1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{q}+1}_{q,1})}\leq C\left(\|u^{n}_{0}\|_{\dot{B}^{\frac{d}{q}-1}_{q,1}}+\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}-1}_{q,1})}\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{q}+1}_{{q},1})}+T\|\tau\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}}_{{q},1})}\right).$ (3.18)
@@
$\|\tau\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}}_{{q},1})}\leq\exp\left(C\|u\|_{L_{T}^{1}(\dot{B}_{{q},1}^{\frac{d}{{q}}+1})}\right)\left(\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{q}}_{{q},1}}+C\|\tau\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}}_{{q},1})}\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{q}+1}_{{q},1})}\right).$ (3.19)
@@
$X_{T}:=\|u(t,\cdot)\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}-1}_{{q},1})}+\|u(t,\cdot)\|_{L^{1}_{T}(\dot{B}^{\frac{d}{q}+1}_{{q},1})},\quad Y_{T}:=\|\tau(t,\cdot)\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}}_{{q},1})}\quad\text{and}\quad Z_{T}:=X_{T}+2CTY_{T}.$
@@
$Z_{T}$ $\leq\left(C+Ce^{CX_{T}}\right)\left(\|u^{n}_{0}\|_{\dot{B}^{\frac{d}{q}-1}_{{q},1}}+T\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{q}}_{{q},1}}+(Z_{T})^{2}\right)\leq Ce^{CZ_{T}}\left(\Gamma_{n}+(Z_{T})^{2}\right).$ (3.20)
@@
$\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}-1}_{{q},1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{q}+1}_{{q},1})}\leq Z_{T}\leq 2C\Gamma_{n}.$ (3.21)
@@
$\|\tau\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}}_{{q},1})}\leq 2\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{q}}_{{q},1}}\leq C\Gamma_{n}\ln n.$ (3.22)
@@
$\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{{q},1}^{\frac{d}{q}+1})}$ $\leq\exp\left(C\|u\|_{L_{T}^{1}(\dot{B}_{{q},1}^{\frac{d}{{q}}+1})}\right)\left(\left\|\tau^{n}_{0}\right\|_{\dot{B}_{{q},1}^{\frac{d}{q}+1}}+C\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{q,1}^{\frac{d}{q}})}\|u\|_{L_{T}^{1}(\dot{B}_{{q},1}^{\frac{d}{{q}}+2})}\right.$
$\quad\left.+C\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{q,1}^{\frac{d}{q}+1})}\|u\|_{L_{T}^{1}(\dot{B}_{q,1}^{\frac{d}{q}+1})}\right)$
$\leq C\Gamma_{n}\left(2^{n}\ln n+\ln n\|u\|_{L_{T}^{1}(\dot{B}_{{q},1}^{\frac{d}{{q}}+2})}+\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{q,1}^{\frac{d}{q}+1})}\right),$
@@
$\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{{q},1}^{\frac{d}{q}+1})}$ $\leq C\Gamma_{n}\ln n\left(2^{n}+\|u\|_{L_{T}^{1}(\dot{B}_{{q},1}^{\frac{d}{{q}}+2})}\right).$ (3.23)
@@
$\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}}_{{q},1})}$ $+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{q}+2}_{{q},1})}\leq\left\|u^{n}_{0}\right\|_{\dot{B}_{q,1}^{\frac{d}{q}}}+C\|u\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{q,1}^{\frac{d}{q}})}\|u\|_{L_{T}^{1}(\dot{B}_{q,1}^{\frac{d}{q}+1})}+T\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{q,1}^{\frac{d}{q}+1})}$
$\leq C\left\{\Gamma_{n}2^{n}+\Gamma_{n}^{-1}2^{-n}+(\Gamma_{n}+\Gamma_{n}^{-1}2^{-2n})\left(\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}}_{{q},1})}+\|u\|_{L_{T}^{1}(\dot{B}_{{q},1}^{\frac{d}{{q}}+2})}\right)\right\},$
@@
$\|u\|_{\tilde{L}^{\infty}_{T}(\dot{B}^{\frac{d}{q}}_{{q},1})}+\|u\|_{L^{1}_{T}(\dot{B}^{\frac{d}{q}+2}_{{q},1})}$ $\leq C\Gamma_{n}2^{n}$
@@
$\|\tau\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{{q},1}^{\frac{d}{q}+1})}$ $\leq C\Gamma_{n}\ln n2^{n}.$
@@ 3.4 Discontinuity of the solution map
Now, we give the lower bound estimation of $\|\tau(T)\|_{\dot{B}^{\frac{d}{p}}_{p,r}}$ with $T=2^{-2n}(\ln n\Gamma_{n}^{2})^{-1}$ which is crucial for the proof of the discontinuity of the solution map.
@@
$\|\tau(t)-\tau^{n}_{0}\|_{\dot{B}^{d}_{{1},1}}$ $\leq Ct\|\tau\|_{\tilde{L}_{t}^{\infty}(\dot{B}^{d}_{{1},1})}\|u\|_{\tilde{L}_{t}^{\infty}(\dot{B}^{d+1}_{{1},1})}+Ct\|\tau\|_{\tilde{L}_{t}^{\infty}(\dot{B}^{d+1}_{{1},1})}\|u\|_{\tilde{L}_{t}^{\infty}(\dot{B}^{d}_{{1},1})}$
@@
$\|\omega(t)-\omega^{n}_{0}\|_{\dot{B}^{d}_{{1},1}}$ $\leq C\|u(t)-u^{n}_{0}\|_{\dot{B}^{d+1}_{{1},1}}$
$\leq C\|u\|_{\tilde{L}_{t}^{\infty}(\dot{B}^{d}_{{1},1})}\|u\|_{L_{t}^{1}(\dot{B}^{d+2}_{{1},1})}+Ct\|u\|_{\tilde{L}_{t}^{\infty}(\dot{B}^{d+3}_{1,1})}+Ct\|\tau\|_{\tilde{L}_{t}^{\infty}(\dot{B}^{d+2}_{{1},1})}$
$\leq C\Gamma^{2}_{n}2^{2n}+Ct\Gamma_{n}2^{4n}+Ct\Gamma_{n}\ln n2^{2n}$
$\leq C\Gamma^{2}_{n}2^{2n}+C(\ln n\Gamma_{n})^{-1}2^{2n},$
@@
$\|P(t)-P_{0}\|_{L^{\infty}}$ $\leq C\|\omega(t)-\omega^{n}_{0}\|_{\dot{B}^{d}_{{1},1}}\|\tau(t)\|_{\tilde{L}_{t}^{\infty}(\dot{B}_{1,1}^{d})}+C\|\tau(t)-\tau^{n}_{0}\|_{\dot{B}^{d}_{1,1}}\|u^{n}_{0}\|_{\tilde{L}_{t}^{\infty}(\dot{B}_{1,1}^{d+1})}$
@@
$\leq$ $~{}CT^{2}\left(\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}+1}_{p,1}}\|u^{n}_{0}\|_{\dot{B}^{\frac{d}{p}+1}_{p,1}}+\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{p,1}}\|u^{n}_{0}\|_{\dot{B}^{\frac{d}{p}+2}_{p,1}}\right)\|u\|_{\tilde{L}_{T}^{\infty}(\dot{B}_{1,1}^{d})}$
@@
$T\|\dot{\Delta}_{n}P_{0}\|_{L^{\infty}}\geq\tilde{c}T2^{2n}\ln n\Gamma_{n}^{2}=\tilde{c}\quad\text{and}\quad\|\tau^{n}_{0}\|_{\dot{B}^{\frac{d}{p}}_{p,r}}\leq C\Gamma_{n}.$ (3.26)